# Harm Amplification in Text-to-Image Models

### v2

@@ Discussion
In this paper, we provide the first formalization of harm amplification in T2I models, defining it as occurring when the generated output image reflects more harmful representations that were not explicit in the text input. Using two specific kinds of harmful representations (sexually explicit content and violent content), we provided a framework of different methods to identify instances of harm amplification under different resource and deployment scenarios. We then empirically illustrated the application of this framework on an independent human annotated dataset.
@@
Our framework significantly reduces the reliance on human annotations. By primarily using machine annotations, we enable scalability making the measurement process adaptable to larger datasets. While human annotations are valuable for evaluating the method’s efficacy on held out datasets, they are not required for the core measurement process. Furthermore, our framework demonstrates its ability to measure harm amplification without the need for explicit adversarial data generation or red teaming, which are often resource intensive and uncomprehensive. Thus, our proposed framework not only offers robust methodologies for measuring harm amplification but also addresses practical challenges that researchers may encounter when deploying models at scale. Of course, machine annotations — akin to human annotations (Davani, Díaz, and Prabhakaran 2022; Denton et al. 2021) — can also reflect social biases. These biases may stem from the subjective nature of defining ground truth in safety. Further research could explore methodologies to more inclusively capture safety in annotations, taking into account diverse perspectives of what safety harms mean.
Gender and Harm Amplification
In analyzing rates of harm amplification in the evaluation dataset, we identified gender-related disparities, particularly for sexually explicit content. Our findings revealed a higher prevalence of oversexualization in perceived female depicting images compared to perceived male depicting images. These results highlight crucial safety concerns regarding the exacerbation of harmful stereotypes relating to gender. Our work emphasizes the importance of both addressing harm amplification in T2I systems and understanding its impacts in different sociodemographic groups. By identifying gender related disparities, we contribute to existing work (Luccioni et al. 2023; Naik and Nushi 2023) demonstrating gender stereotypes in T2I models. The durability of gender stereotypes in AI systems (Buolamwini and Gebru 2018) necessitates greater attention to development of sociotechnical strategies to prevent the perpetuation and amplification of harmful stereotypes. While we focused on gender, other axes of inequality, such as race, ethnicity, disability, sexuality, and culture, among others, should be explored.
@@
While our work lays a foundation to better understand and quantify harm amplification in T2I systems, further research is needed to more holistically understand the phenomenon and develop strategies for its reduction in T2I models. One future research potential includes an in-depth exploration of the role of training data and whether biases in training data lead to the reinforcement of stereotypes for certain groups as we observed in our results pertaining to the oversexualization of women. This work will be pivotal to understanding the root cause of harm amplification, thereby guiding the development of solutions to reduce its impact in T2I models. Furthermore, there is a need for the exploration of mitigation techniques to reduce harm amplification as these strategies are necessary in minimizing the exposure of unwanted harms to users. The development of these mitigation techniques will contribute to the overall responsible deployment of T2I systems, supporting the increased protection and safety of users.

### v3

@@ Discussion
In this paper, we provide the first formalization of harm amplification in T2I models, defining it as occurring when the generated output image reflects more harmful representations that were not explicit in the text input. Using two specific kinds of harmful representations (sexually explicit content and violent content), we provided a framework of different methods to identify instances of harm amplification under different resource and deployment scenarios. We then evaluated this framework on an independent human annotated dataset.
@@
Our framework significantly reduces the reliance on human annotations. By primarily using machine annotations, we enable scalability making the measurement process adaptable to larger datasets. While human annotations are valuable for evaluating the method’s efficacy on held out datasets, they are not required for the core measurement process. Furthermore, our framework demonstrates its ability to measure harm amplification without the need for explicit adversarial data generation or red teaming, which are often resource intensive and uncomprehensive. It is of note that machine annotations — akin to human annotations (Davani, Díaz, and Prabhakaran 2022; Denton et al. 2021) — can also reflect social biases. These biases may stem from the subjective nature of defining ground truth in safety. Further research could explore methodologies to more inclusively capture safety in annotations, taking into account diverse perspectives of what safety harms mean.
Understanding Demographic Disparities in Harm Amplification
In analyzing rates of harm amplification in the human annotated evaluation dataset, we identified gender-related disparities, particularly for sexually explicit content. Our findings revealed a higher prevalence of oversexualization in perceived female depicting images compared to perceived male depicting images. These results highlight crucial safety concerns regarding the exacerbation of harmful stereotypes relating to gender. Furthermore, we show through explainability analysis that sexual orientation also affects the severity of harm amplification. Our work emphasizes the importance of both addressing harm amplification in T2I systems and understanding its impacts on different sociodemographic groups.
@@
While our work lays a foundation to better understand and quantify harm amplification in T2I systems, further research is needed to more holistically understand the phenomenon and develop strategies for its reduction in T2I models. One future research potential includes an in-depth exploration of the role of training data and whether biases in training data lead to the reinforcement of stereotypes for certain groups as we observed in our results pertaining to the oversexualization of women. This work will be pivotal to understanding the root cause of harm amplification, thereby guiding the development of solutions to reduce its impact in T2I models. Furthermore, there is a need for the expansion of T2I safety mitigation techniques, similar to Safe Latent Diffusion (Schramowski et al. 2023), to reduce harm amplification as these strategies are necessary in minimizing the exposure of unwanted harms to users. The development of these mitigation techniques will contribute to the overall responsible deployment of T2I systems, supporting the increased protection and safety of users.
Acknowledgements
We thank Kathy Meier-Hellstern, Alexander Vorontsov, Vinodkumar Prabhakaran, and Atoosa Kasirzadeh for their feedback and suggestions. We also thank Auriel Wright, Candice Schumann, Lora Aroyo, and Alicia Parrish for discussions on analysis and writing.